Circular No. 8515

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005bp IN UGC 10732 Further to IAUC 8509, T. Puckett and A. Pelloni report the discovery

of an apparent supernova (mag 18.8) on an unfiltered CCD frame taken with the 0.60-m automated supernova patrol telescope on Apr. 17.36 UT. The new object was confirmed at mag 18.3 on CCD frames taken on Apr. 18.28. SN 2005bp is located at α = 17^{h}09^{m}43^{s}.22, δ = +36^{o}25′08′′.1 (equinox 2000.0), which is 7′′.6 east and 13′′.7 north of the center of UGC 10732. Nothing is present at this location on images taken by Puckett on 2002 Apr. 16 and 2003 Apr. 1 (limiting mag ∼ 20.0).

SUPERNOVA 2005bq IN IC 4367 Further to IAUC 8507, L. A. G. Monard reports his discovery of an

apparent supernova (mag ∼ 18.3 ± 0.5) on unfiltered CCD images taken with a 0.30-m reflector on Apr. 17.038 UT. The new object, which was confirmed at mag ∼ 18.4 ± 0.4 on images taken on Apr. 17.962 and 18.810, is located at α = 14^{h}05^{m}37^{s}.05, δ = −39^{o}12′04′′.4 (equinox 2000.0), which is 6′′ east and 7′′ north of the nucleus of IC 4367. Nothing is visible at this position on the Digitized Sky Survey (limiting red mag 20.5) or on an image taken on Mar. 5.945 by Monard (limiting red mag 18.5).

SUPERNOVA 2005bl IN NGC 4070 Puckett and A. Langoussis also report an independent discovery of SN

2005bl at mag 17.9 on two images taken on Apr. 18.05 UT, but reporting it as being 13′′.9 east and 11′′.2 south of the center of NGC 4070 (not NGC 4059, as incorrectly given on IAUC 8512 and 8514, due to its erroneous listing at the NED website). Puckett provides the following precise position for SN 2005bl: α = 12^{h}04^{m}12^{s}.32, δ = +20^{o}24′24′′.8 (equinox 2000.0). Nothing was visible at this location on Puckett’s images from 2002 Feb. 1 and 2003 Mar. 8 (limiting mag 19.7–20.0).

COMET C/2000 V4 (SOHO) R. Kracht notes that this tailed comet (which was at mag 7.2 on Nov.

10.896 UT) appears related to C/2001 T5 (continuation to IAUC 8500):

Comet 2000 UT α_{2000} δ_{2000} Inst. F MPEC
C/2000 V4 Nov. 10.772 15^{h}01.^{m}1 −13^{o}40′ C3 RK 2005-G93

2005 April 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green